# TrustEMG-Net: Using Representation-Masking Transformer with U-Net for Surface Electromyography Enhancement

### v1

@@ 1 Introduction
Surface electromyography (sEMG) is a bio-signal that monitors human muscle activity noninvasively. During muscle contraction, the electric potential changes generated by motor units are captured by electrodes placed on the skin. The non-invasive nature makes sEMG widely adopted in various applications, such as neuromuscular system investigation [1], respiration monitoring [2], rehabilitation [3], human–computer interfaces [4], and speech aids [5]. However, its non-invasive measurement also renders sEMG susceptible to various contaminants, such as baseline wander (BW), powerline interference (PLI), electrocardiogram (ECG), motion artifacts (MOA), and white Gaussian noise (WGN) [6, 7]. These contaminants can distort the amplitude and frequency spectrum of the sEMG, hindering information extraction and interpretation [7]. Hence, developing contaminant removal methods for sEMG is crucial to maintaining the efficacy and robustness of sEMG applications.
@@
Recently, neural networks (NNs) have gained prominence for signal enhancement, such as in acoustics [17, 18], ECG [19, 20], and EEG [21]. In these studies, NN-based noise-removal techniques outperform traditional methods, which is primarily attributed to the effective nonlinear mapping capability and data-driven nature of NNs. Various NN models have been incorporated in these studies, including multilayer perceptrons (MLP) [17, 21], convolutional neural networks (CNNs) [21, 19], long short-term memory models (LSTM) [21], fully convolutional neural networks (FCNs) [18, 19], U-Net [20, 22], and Transformer [23]. Several studies have explored NN-based approaches for sEMG noise removal [24, 25]. However, these studies primarily developed techniques using typical MLP-based models [24, 26], or addressed specific noise types (e.g., PLI [24] or ECG [25]). These limitations may constrain the effectiveness and usability of NN-based sEMG denoising. Hence, a more effective and suitable NN model structure for sEMG denoising must be developed, and its robustness and generalizability should be validated for various contaminant types.
To address the limitations of the existing sEMG denoising methods, we propose an NN-based approach called TrustEMG-Net. TrustEMG-Net integrates U-Net with a Transformer encoder layer as a denoising autoencoder (DAE). Previous studies [27] demonstrated that U-Net captures local information excellently through its convolutional layers and residual architecture. Other studies [27, 28] have shown that Transformers are highly effective in modeling global (long-term) information in sequential signals using a self-attention mechanism. Integration automatically utilizes local and global information within the data to remove contaminants. Furthermore, previous studies [29, 30] have confirmed that signals that may combine multiple ingredients can be more easily disentangled and analyzed separately in a high-dimensional representation space compared with the original physical domain. Based on this concept, in this study, we employed the representation-masking (RM) approach for the Transformer encoder to explicitly conduct denoising in a high-level representation space. The experimental results demonstrated that TrustEMG-Net outperforms existing sEMG denoising methods by achieving superior signal quality and low feature extraction errors across various signal-to-noise ratios (SNRs) and contaminant types. The RM approach can also enhance the denoising performance, particularly for narrow-band contaminants. These findings suggest that TrustEMG-Net offers an effective and robust model architecture suitable for sEMG waveform denoising, demonstrating its potential for providing high-quality signals for sEMG applications.
@@ 3.1.1 Clean sEMG database
This study utilized the Non-Invasive Adaptive Prosthetics (Ninapro) database as a clean sEMG data source [42]. Twelve sEMG channels were recorded using active wireless electrodes placed on the upper arm, with a sampling rate of 2 kHz. The Ninapro database contains three subsets of data: DB1, DB2, and DB3. In this study, we adopted DB2, which comprises sEMG data collected from 40 intact subjects and thus has the most extensive subject pool. DB2 has three sessions: Exercises 1, 2, and 3. This study used the first two sessions, in which the subjects performed 17 and 22 types of movements, respectively. Each movement type was executed six consecutive times, with a duration of five seconds, followed by a three-second rest period. Notably, previous studies have viewed this database as a reliable source of clean sEMG data after applying appropriate filtering techniques [43, 25, 44]. We omitted Ninapro DB1 and DB3 as they are unsuitable for clean sEMG datasets. DB1 only provides the root-mean-square rectified version of the sEMG signals and not the raw sEMG signals [42]. As our study focused on denoising raw sEMG signals, DB1 was inappropriate for our purposes. DB3 includes data from only 11 trans-radial amputees, which is a relatively small and unique subject pool. Moreover, DB3 potentially contains more contamination and may not be a reliable source of clean sEMG data [43].
We initiated a series of preprocessing steps to prepare the clean sEMG data for the analysis. The sEMG data were passed through a fourth-order Butterworth bandpass filter with cutoff frequencies of 20 Hz and 500 Hz. Subsequently, the sEMG signals were downsampled to 1 kHz and normalized by dividing them by the maximum absolute value. Finally, we segmented the sEMG signals into 2-s segments and removed the silent ones. Note that although the preprocessing discards sEMG components below 20 Hz and above 500 Hz, this study aimed to restore the main frequency components of sEMG within the 20–500 Hz frequency band. This range contains the most useful information and is commonly employed in various sEMG applications [45, 46, 47].
@@ 3.3 Proposed method
Furthermore, the Transformer encoder employs an RM approach that helps the model learn to denoise at a high-level representation. Masking is a denoising technique often used with generic transformations, such as the short-time Fourier transform [60, 61]. A mask can suppress unwanted signals and preserve the target signals by assigning weights close to zero or one to different parts of a representation. Moreover, many studies have shown that the use of learnable transformations (i.e., NN-based encoders and decoders) with masking techniques can further enhance performance in denoising and signal separation [62, 63]. Generic transformations may not be optimal for various denoising tasks. In TrustEMG-Net, the U-Net’s encoder and decoder serve as learnable transformations; thus, the Transformer encoder can function as a mask predictor to implement the RM technique.
@@ 3.5 Evaluation criteria
The second group of metrics, the feature extraction error, includes the RMSE of the average rectified value (ARV) and mean frequency (MF) feature vectors extracted from noisy and enhanced sEMG. Both the ARV and MF are often used in sEMG applications for clinical evaluation and diagnosis, including trunk muscle fatigue assessment [65, 66], muscle force evaluation [66], and muscle compensation detection in rehabilitation [67]. The ARV is calculated as follows:
@@ 4 Results
This study analyzed the experimental results from three perspectives: overall performance, performance under different SNR inputs, and performance with different types of contaminants. In the first subsection, the proposed method is compared with five existing sEMG denoising algorithms: IIR filters [9, 13, 6], TS with IIR filters [31, 69, 9], and EMD-based [13, 70], CEEMDAN-based [35, 11, 71], and VMD-based sEMG denoising methods [13, 36]. The second subsection describes the ablation study and compares the proposed method with several CNN-based architectures (CNN, FCN, and U-Net) because convolutional layers are suitable for processing waveform input data. Moreover, a direct mapping (DM) approach for the Transformer encoder was implemented and denoted as TrustEMG-Net(DM) for comparison. The difference between the DM and RM is shown in Fig. 2. A detailed implementation of the compared methods and NN models is available in our supplementary material and on the GitHub page [^2] https://github.com/eric-wang135/TrustEMG/tree/main/doc.
@@ 5.1 Comparison with existing sEMG denoising methods
Figure 6: Investigation into the importance of preserving and suppressing specific features in the representation. The observations highlight the importance of optimizing the mask to effectively suppress contaminants while preserving essential sEMG features for optimal denoising results. (a) The input sEMG signal was contaminated by PLI (a sinusoidal wave of about 60 Hz) at an SNR of -14 dB. (b) We manually adjusted the weights of the original predicted mask and observed changes in the output waveform. Two modifications are applied: increasing the weights of low channels to 1 (preserving the initially suppressed features) and decreasing the weights of high channels to 0 (suppressing the initially preserved features). (c) As the mask weights increased, the signal quality deteriorated owing to the gradual emergence of PLI, indicating that the features initially suppressed by the mask are closely associated with contaminants. Conversely, (d) as the mask weights increased, the signal quality decreased owing to sEMG distortion, suggesting that preserved features are more related to clean sEMG signals.
@@
Despite the superiority of TrustEMG-Net across various contaminant types, we observed that existing sEMG denoising methods, including simple IIR filters, outperformed TrustEMG-Net under BW contamination, as shown in Table 2. This outcome can be attributed to the characteristics of BW, which predominantly consists of frequency components below 10 Hz. These frequencies do not overlap with the primary sEMG frequency range (20–500 Hz). Consequently, unlike the other contaminant types evaluated in this study, BW can be effectively removed without distorting the sEMG using conventional methods. This observation suggests that NN-based methods may not be essential for BW removal when targeting the sEMG frequency range of 20–500 Hz.
@@ 5.2.2 Representation-masking Transformer
When the mask weights of the low-weight channels were increased to preserve the initially suppressed features, the contaminant (PLI) became apparent in the output waveform, degrading the signal quality. Conversely, reducing the mask weights of the high-weight channels to suppress the initially preserved features resulted in noticeable distortion of the sEMG signal. Therefore, optimal denoising requires a mask to selectively suppress and preserve specific features.
@@ 5.3 Computational complexity analysis
Table 7 lists the computational requirements of the baseline and NN-based sEMG denoising methods. The computation time for the baseline methods is specified for the CPU, whereas the NN-based methods include model parameters, floating-point operations (FLOPs), and GPU and CPU computation times. The computation time indicates the duration for processing a two-second noisy sEMG segment contaminated with a five-type compound contaminant on the Nvidia GeForce RTX 3090 GPU and Intel Xeon Silver 4416+ CPU.
Compared with decomposition-based sEMG denoising methods (EMD, CEEMDAN, and VMD), TrustEMG-Net demonstrates superior efficiency by requiring less computation time, in addition to its effectiveness. This difference largely results from the iterative computational or optimization processes involved in decomposition algorithms. Among NN-based methods, a trade-off exists between the denoising performance and computational efficiency. TrustEMG-Net achieved better denoising results than FCN and U-Net at the cost of more parameters, FLOPs, and computation time. TrustEMG-Net(DM) shares the same parameter count as TrustEMG-Net with a slight increase in computation time owing to the RM technique. Future research could explore techniques such as parameter pruning and quantization [75] to mitigate the computational demands for resource-limited or time-sensitive sEMG applications.
@@ Acknowledgment
We thank Professor Juliano Machado from the Department of Basic, Technical, and Technological Education at Sul-Rio-Grandense Federal Institute for generously providing us with the MOA dataset used in this research. This work was supported by the Academia Sinica under Grant ASGC-111-M01.

### v2

@@ 1 Introduction
Surface electromyography (sEMG) is a bio-signal that monitors human muscle activity. During muscle contraction, the electric potential changes generated by motor units are captured by electrodes placed on the skin. The non-invasive nature makes sEMG widely adopted in various applications, such as neuromuscular system investigation [1], respiration monitoring [2], rehabilitation [3], human–computer interfaces [4], and speech aids [5]. However, its non-invasive measurement also renders sEMG susceptible to various contaminants, such as baseline wander (BW), powerline interference (PLI), electrocardiogram (ECG), motion artifacts (MOA), and white Gaussian noise (WGN) [6, 7]. These contaminants distort the amplitude and frequency spectrum of the sEMG, hindering information extraction and interpretation [7]. Hence, developing contaminant removal methods for sEMG is crucial to maintaining the efficacy and robustness of sEMG applications.
@@
Recently, neural networks (NNs) have gained prominence for signal enhancement, such as in acoustics [17, 18], ECG [19, 20], and EEG [21]. In these studies, NN-based noise-removal techniques outperform traditional methods, which is primarily attributed to the effective nonlinear mapping capability and data-driven nature of NNs. Various NN models have been incorporated in these studies, including multilayer perceptrons (MLP) [17, 21], convolutional neural networks (CNNs) [21, 19], long short-term memory models (LSTM) [21, 22], fully convolutional neural networks (FCNs) [18, 19], U-Net [20, 23], and Transformer [24]. Several studies have explored NN-based approaches for sEMG noise removal [25, 26]. However, these studies primarily developed techniques using typical MLP-based models [25, 27], or addressed specific noise types (e.g., PLI [25] or ECG [26]). These limitations may constrain the effectiveness and usability of NN-based sEMG denoising. Hence, a more effective and suitable NN model structure for sEMG denoising must be developed, and its robustness and generalizability should be validated for various contaminant types.
To address the limitations of the existing sEMG denoising methods, we propose an NN-based approach called TrustEMG-Net. TrustEMG-Net integrates U-Net with a Transformer encoder layer as a denoising autoencoder (DAE). Previous studies [28, 29] demonstrated that U-Net captures local information excellently through its convolutional layers and residual architecture. Other studies [29, 30] have shown that Transformers are highly effective in modeling global (long-term) information in sequential signals using a self-attention mechanism. Integration automatically utilizes local and global information within the data to remove contaminants. Furthermore, previous studies [31, 32, 33] have confirmed that signals that may combine multiple ingredients can be more easily disentangled and analyzed separately in a high-dimensional representation space compared with the original physical domain. Based on this concept, in this study, we employed the representation-masking (RM) approach for the Transformer encoder to explicitly conduct denoising in a high-level representation space. The experimental results demonstrated that TrustEMG-Net outperforms existing sEMG denoising methods by achieving superior signal quality and low feature extraction errors across various signal-to-noise ratios (SNRs) and contaminant types. The RM approach can also enhance the denoising performance, particularly for narrow-band contaminants. These findings suggest that TrustEMG-Net offers an effective and robust model architecture suitable for sEMG waveform denoising, demonstrating its potential for providing high-quality signals for sEMG applications.
@@ 3.1.1 Clean sEMG database
This study utilized the Non-Invasive Adaptive Prosthetics (Ninapro) database as a clean sEMG data source [45]. Twelve sEMG channels were recorded using active wireless electrodes placed on the upper arm, with a sampling rate of 2 kHz. The Ninapro database contains three subsets of data: DB1, DB2, and DB3. In this study, we adopted DB2, which comprises sEMG data collected from 40 intact subjects and thus has the most extensive subject pool. DB2 has three sessions: Exercises 1, 2, and 3. This study used the first two sessions, in which the subjects performed 17 and 22 types of movements, respectively. Each movement type was executed six consecutive times, with a duration of five seconds, followed by a three-second rest period. Notably, previous studies have viewed this database as a reliable source of clean sEMG data after applying appropriate filtering techniques [46, 26, 47]. We omitted Ninapro DB1 and DB3 as they are unsuitable for clean sEMG datasets. DB1 only provides the rectified version of the sEMG signals and not the raw sEMG signals [45]. As our study focused on denoising raw sEMG signals, DB1 was inappropriate for our purposes. DB3 includes data from 11 transradial amputees, which is a relatively small and unique subject pool. Moreover, DB3 potentially contains more contamination and may not be a reliable source of clean sEMG data [46].
We initiated a series of preprocessing steps to prepare the clean sEMG data. The sEMG data were passed through a fourth-order Butterworth bandpass filter with cutoff frequencies of 20 Hz and 500 Hz. Subsequently, the sEMG signals were downsampled to 1 kHz and normalized by dividing them by the maximum absolute value. Finally, we segmented the sEMG signals into 2-s segments and removed the silent ones. Note that although the preprocessing discards sEMG components below 20 Hz and above 500 Hz, this study aimed to restore the main frequency components of sEMG within the 20–500 Hz frequency band. This range contains the most useful information and is commonly employed in various sEMG applications [48, 49, 50].
@@ 3.3 Proposed method
Furthermore, the Transformer encoder employs an RM approach that helps the model learn to denoise at a high-level representation. Masking is a denoising technique often used with generic transformations, such as the short-time Fourier transform [63, 64]. A mask can suppress unwanted signals and preserve the target signals by assigning weights close to zero or one to different parts of a representation. Moreover, many studies have shown that the use of learnable transformations (i.e., NN-based encoders and decoders) with masking techniques can further enhance performance in denoising and signal separation [65, 66, 67]. Generic transformations may not be optimal for various denoising tasks. In TrustEMG-Net, the U-Net’s encoder and decoder serve as learnable transformations; thus, the Transformer encoder can function as a mask predictor to implement the RM technique.
@@ 3.5 Evaluation criteria
The second group of metrics, the feature extraction error, includes the RMSE of the average rectified value (ARV) and mean frequency (MF) feature vectors extracted from noisy and enhanced sEMG. Both the ARV and MF are often used in sEMG applications for clinical evaluation and diagnosis, including trunk muscle fatigue assessment [69, 70], muscle force evaluation [70, 71], and muscle compensation detection in rehabilitation [72]. The ARV is calculated as follows:
@@ 4 Results
This study analyzed the experimental results from three perspectives: overall performance, performance under different SNR inputs, and performance with different types of contaminants. In the first subsection, the proposed method is compared with five existing sEMG denoising algorithms: IIR filters [9, 13, 6], TS with IIR filters [34, 74, 9], and EMD-based [13, 75], CEEMDAN-based [38, 11, 76], and VMD-based sEMG denoising methods [13, 39]. The second subsection describes the ablation study and compares the proposed method with several CNN-based architectures (CNN, FCN, and U-Net) because convolutional layers are suitable for processing waveform input data. Moreover, a direct mapping (DM) approach for the Transformer encoder was implemented and denoted as TrustEMG-Net(DM) for comparison. The difference between the DM and RM is shown in Fig. 2. A detailed implementation of the compared methods and NN models is available on the GitHub page [^2] https://github.com/eric-wang135/TrustEMG/tree/main/doc.
@@ 5.1 Comparison with existing sEMG denoising methods
Figure 6: Investigation into the importance of preserving and suppressing specific features in the representation. The observations highlight the importance of optimizing the mask to effectively suppress contaminants while preserving essential sEMG features for optimal denoising results. (a) The input sEMG signal was contaminated by PLI at an SNR of -14 dB. (b) We manually adjusted the weights of the original predicted mask and observed changes in the output waveform. Two modifications are applied: increasing the weights of low channels to 1 (preserving the initially suppressed features) and decreasing the weights of high channels to 0 (suppressing the initially preserved features). (c) As the mask weights increased, the signal quality deteriorated owing to the gradual emergence of PLI, indicating that the features initially suppressed by the mask are closely associated with contaminants. Conversely, (d) as the mask weights increased, the signal quality decreased owing to sEMG distortion, suggesting that preserved features are more related to clean sEMG signals.
@@
Despite the superiority of TrustEMG-Net across various contaminant types, we observed that existing sEMG denoising methods, including simple IIR filters, outperformed TrustEMG-Net under BW contamination, as shown in Table. 2. This outcome can be attributed to the characteristics of BW, which predominantly consists of frequency components below 10 Hz. These frequencies do not overlap with the primary sEMG frequency range (20–500 Hz). Consequently, unlike the other contaminant types evaluated in this study, BW can be effectively removed without distorting the sEMG using conventional methods. This observation suggests that NN-based methods may not be essential for BW removal when targeting the sEMG frequency range of 20–500 Hz.
@@ 5.2.2 Representation-masking Transformer
When the mask weights of the low-weight channels were increased to preserve the initially suppressed features, the contaminant (PLI) became apparent in the output waveform, degrading the signal quality. Conversely, reducing the mask weights of the high-weight channels to suppress the initially preserved features resulted in noticeable distortion of the sEMG signal. Hence, optimal denoising requires a mask to selectively suppress and preserve specific features.
@@ 5.3 Computational complexity analysis
Table 7 lists the computational requirements of the baseline and NN-based sEMG denoising methods. The computation time for the baseline methods is specified for the CPU, whereas the NN-based methods include model parameters, floating-point operations (FLOPs), and GPU and CPU computation times. The computation time indicates the duration for processing a 2-s noisy sEMG segment contaminated with a five-type compound contaminant on the Nvidia GeForce RTX 3090 GPU and Intel Xeon Silver 4416+ CPU.
Compared with decomposition-based sEMG denoising methods (EMD, CEEMDAN, and VMD), TrustEMG-Net demonstrates superior efficiency by requiring less computation time, in addition to its effectiveness. This difference largely results from the iterative computational or optimization processes involved in decomposition algorithms. Among NN-based methods, a trade-off exists between the denoising performance and computational efficiency. TrustEMG-Net achieved better denoising results than FCN and U-Net at the cost of more parameters, FLOPs, and computation time. TrustEMG-Net(DM) shares the same parameter count as TrustEMG-Net with a slight increase in computation time owing to the RM technique. Future research could explore techniques such as parameter pruning and quantization [80, 81], or more efficient sequence-to-sequence models [82] to mitigate the computational demands for resource-limited or time-sensitive sEMG applications.
@@ Acknowledgment
We thank Professor Juliano Machado from the Department of Basic, Technical, and Technological Education at Sul-Rio-Grandense Federal Institute for providing us with the MOA dataset used in this research. This work was supported by the Academia Sinica under Grant ASGC-111-M01.